# Exploring wetting dynamics of water nanodroplet on nano-textured surfaces: atomistic insights

### v1

@@ Abstract
The static and dynamic wetting behaviors of a water nanodroplet are investigated on flat and structured titanium (Ti) surfaces by using a fully atomistic simulation. Firstly, the results obtained for a water nanodroplet on a flat Ti surface agree with the static contact angle reported in the literature. Furthermore, to unveil the mechanisms of the water nanodroplet behavior on nano-textured surfaces, we analyze the impact of the roughness factors, the structure period, and sizes compared to the nanodroplet diameter. The introduction of the surface roughness produces an inhomogeneous spreading of the nanodroplet. The degree of roughness and the specific dimensions of the surface grooves determine these wetting processes. Surfaces with larger grooves and spacing tend to support more extensive spreading, while those with narrower grooves and higher roughness limit spreading. The obtained results also demonstrate damping in droplet vibrations accompanied by a gradual decrease in the dynamic contact angle of water nanodroplets on both flat and rough Ti surfaces, indicating a transition to a hydrophobic state. This effect is caused by the atomic bonding of water atoms to the surface ones and is shown to strongly depend on the surface roughness factors. The simulation results provide atomic-level insights into the wetting mechanisms on surfaces featuring nanometric and sub-micrometric roughness, typically induced by femtosecond laser processing.
Keywords: wetting, contact angle, titanium, femtosecond laser, roughness, molecular dynamics, nanodroplet
@@ 1 Introduction
Ti and its alloys are highly promising materials, especially for biomedical applications due to their biocompatibility, high strength, and fracture toughness [1]. These characteristics make them ideal for dental and orthopedic implants. It is well known that one of the most effective ways to improve Ti bone integration with these implants is by introducing appropriate surface roughness [2]. Surface texturing may also improve its antibacterial properties. Among several methods of surface treatment, femtosecond laser processing has notable advantages. These include high processing resolution and a minimal heat-affected zone [3]. Additionally, this method allows one to fabricate complex 2D and 3D structures [4, 5]. In this context, adjusting the wettability of the surface plays a fundamental role, as it dictates how the implant interacts with surrounding tissues and promotes the cell attachment to prevent bacterial growth [2, 6]. Therefore, a deep understanding of these wetting properties is very important.
Several models have been developed to account for surface topography and heterogeneities, taking into account chemical patterns and impurities on the textured surfaces [7]. For surfaces treated with femtosecond lasers, the experimentally observed apparent contact angle (CA) often differs from the predictions of Young’s model [8, 9, 10, 11]. This discrepancy is theoretically explained by using the Wenzel and Cassie-Baxter equations that account for these surface irregularities [12, 13]. The Wenzel model assumes that the liquid fully penetrates the solid’s roughness and fills the gaps establishing a microscopic contact angle [12]. In contrast, the Cassie-Baxter model suggests that the liquid does not infiltrate the grooves. Instead, a significant portion of the liquid-gas interface is pinned to the surface due to air entrapment [13].
These models have been consistently used within the laser community to quantify the wettability of laser-treated surfaces [14]. The Wenzel and Cassie-Baxter equations are typically applied to homogeneous and isotropic surfaces where the nanodroplet size to roughness scale ratio is large. However, it is well known that micro/nanostructures created with femtosecond lasers can exhibit heterogeneity [15]. Such heterogeneous topography can complicate the application of these models.
In addition, the nature of nano-surface processing and the challenges of conducting experiments at the nanoscopic scale remain challenging. These challenges restrict the capacity to improve our understanding of the interactions between nanodroplets and these specially textured surfaces. Therefore, Molecular Dynamics (MD) simulations are adequate to investigate the nanodroplet wetting properties on nanometer surface relief as reported in previous studies [16, 17, 18]. This approach is interesting because it allows for an easy assessment of the total energy resulting from interactions between atoms in a system. It also allows us to study droplet stability and its dynamic behavior.
@@
Several studies were already focused on understanding the wettability of titanium dioxide ($TiO_{2}$) using atomistic approaches. For example, Park and Aluru [16] used MD simulations to show that the wettability of a $TiO_{2}$ surface increases with rising temperature due to a decrease in hydrogen bonding and surface tension. Similarly, Ohler and Langel [17] proposed a new model for the CA of water on $TiO_{2}$ surfaces using MD simulations. They argue that the macroscopic CA is not determined by the direct interaction of bulk water with the $TiO_{2}$ surface but rather by an interface between a thin, ordered layer of water molecules on the surface and the bulk water droplet.
@@
Recently, Zhu et al. [18] investigated the wettability of three different phases of $TiO_{2}$ thin films: anatase, brookite, and rutile. The study found that the wettability of the $TiO_{2}$ surface depends on the crystal phase, with anatase being the most wettable (CA of $43.7^{\circ}$) and rutile the least wettable (CA of $73.9^{\circ}$). The authors attribute this difference to variations in the structural connection and arrangement of surface microtopography among the three phases. The study also found that increasing the roughness of the anatase surface decreases its wettability, further highlighting the role of surface morphology in determining wettability.
To the best of our knowledge, the wettability of Ti has not been extensively studied using MD simulations. Given the insights gained from MD studies on $TiO_{2}$, a similar atomistic approach to studying Ti wettability could provide a molecular-level understanding of the fundamental interactions, which can offer new perspectives and potentially lead to increased control over surface properties for various applications. In our recent study using hydrodynamic simulation, we demonstrated that to achieve more hydrophobic behavior on laser-textured surfaces, more attention should be paid to the height of the relief, and caution should be exercised when varying the period of the relief during femtosecond laser processing [19, 20].
In this study, MD is used to investigate the wettability of a water nanodroplet on flat and rough Ti surfaces similar to that of the rippled metasurfaces commonly made by femtosecond laser with structure sizes of a few tens of nanometers. Static and dynamic wetting phenomena are both examined by covering flat and structured surfaces. Furthermore, we measured both the static of water nanodroplets and compared to the prediction of classical wetting models discussed above. By comparing these results with experimental data and the classical wetting models, we provide deep insights into the mechanisms at play allowing us a better understanding of the wettability of the structured Ti surfaces.
@@
width = colspec = —Q[1,c,m]—Q[1,c,m]—Q[1,c,m]—Q[1,c,m]—, hlines, vlines, row1 = font=, rowodd = gray9, Pair $\sigma$ (nm) $\varepsilon$ (kJ/mol) Charge (e)
@@ 2 Simulations model and details
The short-range LJ potential ($U_{non-bond}(r_{ij})$) represents the Van der Waal force between two non-bonding atoms. As shown in Equ. 1, it is defined by a conventional 12-6 function such that $r_{ij}$ is the distance between the two atoms i and j, $\sigma_{ij}$ is the distance at which the interaction energy between two atoms is minimal, and $\varepsilon_{ij}$ is the depth of the potential well. In the Coulombic term ($U_{e}(r_{ij})$), $\epsilon$ represents the permittivity in vacuum, $q_{i}$ and $q_{j}$ denote the charges of the interaction sites. The OH bond length is 1 Å and the angle between H-O-H is fixed at 109.47. To make the water molecule more rigid during the MD simulation, the SHAKE algorithm was used to constrain bond length and angle [23]. Long-range electrostatic interactions were computed using the Particle-Particle Particle-Mesh (PPPM) method [24]. The PPPM is more efficient than the Ewald method [25] because it can handle long-range interactions for periodic systems [26]. Non-bonding interactions, including both Van der Waals and electrostatic forces, were truncated at a cut-off distance of 12 Å. The interactions between Ti and O atoms were modeled using the 12-6 LJ potential, with parameters determined by the Lorentz-Berthelot mixing rule [27], as shown in Equ. 2.
@@
It should be noted that the interactions between Ti and H atoms are neglected in this study. This simplification based on the omission of LJ parameters for H atoms in the SPC/E water model is a well-established practice supported by extensive research and literature. The primary reasons for this simplification are computational efficiency and the focus to accurately capture the critical electrostatic interactions that determine the behavior of the water molecules [28, 29]. Additionally, because the bonds in the SPCE model of water are rigid, the H atoms are restricted from freely rotating and forming independent interactions with Ti [21]. Therefore, the H atoms are primarily involved in H bonding with neighboring water molecules [28], further reducing their availability for direct interaction with Ti. It is worth mentioning that neglecting H-substrate interaction may underestimate the influence of wetting behavior on highly reactive (laser-treated) surfaces that can readily form H bonds with the H atoms of water.
@@
width = colspec = —Q[1,c,m]—Q[1,c,m]—Q[1,c,m]—Q[1,c,m]—Q[1,c,m]—, hlines, vlines, row1 = font=, rowodd = gray9, Surface (Surf) h (nm) w (nm) s (nm) p (nm)
@@
In addition, Ti-Ti interactions were calculated using the LJ potential. The LJ potential parameters for Ti interactions used in this study are modified in such a way that they can replicate the physical properties of Ti. The parameters have been validated against Density Functional Theory (DFT) calculations and experimental data. These parameters accurately reproduce the lattice constant and independent elastic constants, as detailed in the Supporting Information (1.1). The LJ potential parameter values used in the MD simulations of this study are provided in Table 1.
The nanodroplet corresponds to a sphere with a radius of 3.0 $nm$ and contains $3009$ water molecules. Previous studies have consistently demonstrated that droplets comprising 2000 atoms (or more) consistently produce reliable results [30]. Temperature equilibration of the nanodroplet and the Ti substrate was performed separately. The nanodroplet was equilibrated using a time step of 1 $fs$ at 300 $K$ for approximately 1 $ns$ in the NVT ensemble. This duration was sufficient to achieve a stable nanodroplet configuration. To validate this configuration, the radial distribution functions (g(r)) obtained from our simulations were compared with the experimental data [31] and the SPC/E model [21], as shown in Fig. 1. The comparison indicates good agreement between our study, the experimental data, and the SPC/E model with slight deviations observed below the peaks of g(r). This agreement demonstrates that the SPC/E water model used in our MD simulations is appropriate and reliable.
@@
Figure 2: (a) A fence-like nanostructure on Ti surface with heights ($h$), widths ($w$), spacing ($s$), and the periodicity ($p$). (b) Initial atomistic configuration of water nanodroplet on the flat Ti surface.
Figure 3: Snapshots of the side and front views of the spreading process of a water nanodroplet on Ti surfaces with different roughness scales as a function of time.
The flat Ti substrate corresponds to a block of dimension $14.5$$\times$ $19.5$$\times$ $1.4$ nm3 containing $22344$ atoms arranged in a Hexagonal Close-Packed (HCP) structure along [0-110] [10-10] [0001] directions. Subsequently, a nanometric pattern similar to laser-textured ones, having heights ($h$), widths ($w$), spacing ($s$), and a periodicity ($p$) as depicted in Fig. 2 (a), was introduced on Ti surfaces [20, 32, 33]. In our previous studies, it has been established that the structural parameters $h,w,s,$ and $p$ are adjustable based on the topography profile of the laser-patterned surfaces [20]. The adjustable parameters used in the wetting studies are summarised in Table 2. For the sack of clarity and to facilitate the adjustment of these parameters, we use the roughness factor ($r$) and fractional area ($f$) to quantify them, as shown in Equ. 3 below.
@@
The bulk Ti substrate was equilibrated using the NVT ensemble at 300 $K$ for 1 $ns$ with a time step of 1 $fs$. In addition, the equilibrated spherical nanodroplet was positioned on the equilibrated Ti surface, as shown in Fig. 2 (b). In this study, the flat and rough Ti substrates were completely frozen during the MD simulation. The fixed substrate in MD simulations, particularly in wetting studies, is a well-documented practice aimed at a reduced computational cost [28], maintaining numerical stability [34] and ensuring an accurate representation of the physical system [35, 36]. Interestingly, simulations with a fixed surface compared to a flexible surface have shown a negligible effect on the CA [22]. In addition to the simulation set-up, the size of the simulation box was set to $14.5$$\times$ $19.5$$\times$ $15.4$ nm3. Periodic boundary conditions were applied in the x, y, and z directions during the simulation of the water nanodroplet on the Ti surfaces. The entire simulation was carried out for 1 $ns$ in the NVT ensemble with a time step of 1 $fs$. MD calculations were performed using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) software [37]. Visualization and analysis of the water nanodroplet on Ti surfaces were conducted using OVITO software [38].
@@ 3.1 Dynamic wetting behavior of Ti surfaces
To understand the fundamental process of nanodroplet spreading on the flat and nanometric Ti surfaces, the dynamics of the nanodroplet on the corresponding surfaces are examined. Thus, Fig. 3 shows the snapshots of the side and front views of the spreading process of a water nanodroplet on Ti surfaces with different roughness scales as a function of time.. The droplet on the flat surface (Surf-1) is initially spherical at 0 ps, indicating minimal spreading. Between 30 and 50 ps, the droplet starts to spread gradually, with a noticeable decrease in height. From 200 to 1000 ps, further spreading occurs, and the droplet becomes more stable with a large base and minimal height.
@@
In the case of Surf-2, the droplet spreads and begins to interact with the rough surface by 30 ps. As time progresses, between 200 and 1000 ps, the droplet flattens out, forming a wide base that settles into the roughness grooves. This indicates a substantial degree of spreading. In contrast, for Surf-3, the spreading process is restricted by the roughness grooves. The droplet conforms to these grooves rather than spreading freely. This limitation occurs because Surf-3 has a greater number of grooves beneath the droplet compared to Surf-2. The equal distribution of w and s for Surf-3, as opposed to Surf-2, provides evidence for this increased groove density.
@@
Fig. 5 illustrates the dynamic CA of water nanodroplets as a function of time for Ti surfaces with varying roughness levels. At 30 ps, the CAs are generally high, indicating that the droplets initially exhibit minimal spreading on all surfaces. The initial CAs vary significantly with the roughness of the surfaces; for instance, Surf-1 starts at approximately $115^{\circ}$, whereas Surf-4 begins at around $133^{\circ}$. As time progresses, the CAs decrease, signifying that the droplets spread out on the surfaces. This decrease is most rapid during the initial 50 ps, reflecting the dynamic wetting process. Beyond 50 ps, the rate of decrease in CA slows down as the system approaches equilibrium, and the contact angles stabilize. Notably, Surf-5 exhibits a higher CA compared to Surf-4, 3, 2, and 1. This behavior indicates that rougher surfaces, such as Surf-5, prevent wetting more effectively than smoother ones, leading to higher CA. The higher roughness in Surf-5 makes it trap air pockets under the droplet, reducing the actual contact area between the droplet and the surface.
@@
Thus, in general, the CA increases as roughness rises shifting the wettability of Ti surfaces from hydrophilic (Surf-1) to hydrophobic state for other surfaces. These observations for Surf-2, 3, and 4 align with the main assumption of the Wenzel model. The Cassie-Baxter model somewhat better describes Surf-5, where air pockets exist between the droplet and the rough Ti surface, generally leading to higher CAs and less wetting. Interestingly, despite small sizes, the initial rapid decrease in CA and an increase in spreading radius is similar to the dynamic wetting theories and previous simulation studies, where the droplet adjusts its shape quickly upon initial contact to minimize the surface energy. It is important to mention, however, that the nandroplet undergoes significant vibration on all the surfaces at 50 ps to 1000 ps, as seen in Fig. 3, 4 and 5. The energy fluctuations and the resultant variability in surface tension at the nanoscale lead to a set of dynamic instabilities [39]. Additionally, intrinsic thermal motion and molecular interactions within the droplet cause continuous vibrations [40]. The higher surface area-to-volume ratio in nanoscale droplets can increase these effects, leading to noticeable oscillations [39, 40].

### v2

# Atomistic Simulations of Wetting Dynamics of Water Nanodroplets on Nanotextured Titanium: Implications for Medical Implants

@@ Abstract
The development of high-performance biomedical implants requires a deep understanding of the molecular interactions between water molecules and titanium (Ti) surfaces. In this study, fully atomistic molecular dynamics simulations were used to study the static and dynamic wetting behavior of water nanodroplets on both flat and femtosecond laser-induced nanotextured Ti surfaces. Our findings reveal a clear transition from Wenzel to Cassie-Baxter wetting states as surface roughness increases, significantly affecting droplet spreading. We also observe the damping of nanodroplet vibrations and a roughness-dependent shift toward hydrophobicity, driven by stronger atomic interactions between water molecules and surface atoms. Furthermore, the interaction energy between water droplets and nano-textured Ti surfaces decreases with increasing roughness, reinforcing the observed changes in wettability. The discrepancies observed between classical wetting models and nanoscale behavior emphasize the limitations of current theoretical approaches and the importance of developing more advanced models. This study provides valuable insights into optimizing Ti surface properties for improved implant performance through controlled wettability.
Keywords: titanium, wetting, contact angle, femtosecond laser, roughness, molecular dynamics, water, nanodroplet
@@ 1 Introduction
Titanium (Ti) and its alloys hold immense promise, particularly in biomedical applications, due to their exceptional biocompatibility, high strength, and fracture toughness [1]. These properties make them ideal candidates for dental and orthopedic implants. A well-established strategy for improving the integration of Ti implants with bone involves the introduction of appropriate surface roughness [2]. Surface texturing has also been shown to improve antibacterial properties [3]. Among the various surface treatment methods, femtosecond laser processing stands out due to its high processing resolution, minimal heat-affected zone [4], and the ability to fabricate complex 2D and 3D structures [5, 6].
The wettability of laser-treated implant surface plays a pivotal role in influencing its interaction with surrounding tissues, promoting cell attachment, and preventing bacterial growth [2, 7]. A comprehensive understanding of the wetting properties of Ti surfaces is important. While several models have been developed to account for surface topography and heterogeneities [8], the experimentally observed apparent contact angle (CA) on femtosecond laser-treated surfaces often deviates from the predictions of Young’s model [9, 10, 11, 12]. This discrepancy is typically explained using the Wenzel and Cassie-Baxter equations, which consider surface irregularities [13, 14]. However, these models are primarily applicable to homogeneous and isotropic surfaces with large droplet-to-roughness scale ratios, and may not fully capture the variations of heterogeneous micro/nanostructures created by femtosecond lasers [15]. On the other hand, studies have shown that micro/nanostructures fabricated using femtosecond lasers exhibit dynamic wetting properties, where the contact angle of a liquid changes over time due to advancing and receding contact angles, influenced by surface topography and energy [5, 16, 17].
The challenges associated with nano-surface processing and experimentation at the nanoscale have significantly limited our understanding of nanodroplet interactions with textured surfaces. To overcome these limitations, Molecular Dynamics (MD) simulations have emerged as a powerful tool for investigating the wetting properties of nanodroplets on nanometer-scale surface reliefs. This approach holds significant potential for various applications, particularly in understanding how nanostructures created by femtosecond laser texturing affect the attachment of cells to medical implants [18, 19, 20]. By optimizing these surface textures, femtosecond laser technology can promote specific cellular responses, improving the integration of artificial cells with medical implants and advancing the overall success of implant procedures [21, 22].
@@
Previous research has extensively studied the wetting behavior of titanium dioxide (TiO2) using MD simulations. For example, Park and Aluru [23] used MD simulations to show that the wettability of a TiO2 surface increases with rising temperature due to a decrease in hydrogen bonding and surface tension. Similarly, Ohler and Langel [24] proposed a new model for the CA of water on TiO2 surfaces using MD simulations. They argue that the macroscopic CA is not determined by the direct interaction of bulk water with the TiO2 surface but rather by an interface between a thin, ordered layer of water molecules on the surface and the bulk water droplet.
@@
Recently, Zhu et al. [25] investigated the wettability of three different phases of TiO2 thin films: anatase, brookite, and rutile. The study found that the wettability of the TiO2 surface depends on the crystal phase, with anatase being the most wettable (CA of 43.7) and rutile the least wettable (CA of 73.9). The authors attribute this difference to variations in the structural connection and arrangement of surface microtopography among the three phases. The study also found that increasing the roughness of the anatase surface decreases its wettability, further highlighting the role of surface morphology in determining wettability. However, despite the wealth of research on TiO2, the wettability of pure Ti remains largely unexplored.
In this study, we address this gap by employing MD simulations to investigate the wettability of water nanodroplets on Ti, both on flat and nano-textured surfaces, mimicking the rippled metasurfaces commonly produced by femtosecond laser processing with structure sizes of a few tens of nanometers. This research provides new atomic-level insights into how water interacts with Ti, particularly under varying surface roughness. Our study reveals the influence of nanostructuring on the wettability of Ti, which is essential for optimizing surface properties for medical implants. Additionally, by examining both static and dynamic wetting phenomena, we provide a comprehensive understanding of the wetting mechanisms, thus offering data that can bridge the gap between classical models and experimental observations.
@@
width = colspec = |Q[1,c,m]|Q[1,c,m]|Q[1,c,m]|Q[1,c,m]|, hlines, vlines, row1 = font=, rowodd = gray9, Pair $\sigma$ (nm) $\varepsilon$ (kcal/mol) Charge (e)
@@ 2 Simulations model and details
The short-range LJ potential ($U_{non-bond}(r_{ij})$) represents the Van der Waal force between two non-bonding atoms. As shown in Equ. 1, it is defined by a conventional 12-6 function such that $r_{ij}$ is the distance between the two atoms i and j, $\sigma_{ij}$ is the distance at which the interaction energy between two atoms is minimal, and $\varepsilon_{ij}$ is the depth of the potential well. In the Coulombic term ($U_{e}(r_{ij})$), $\epsilon$ represents the permittivity in vacuum, $q_{i}$ and $q_{j}$ denote the charges of the interaction sites. The OH bond length is 1 Å and the angle between H-O-H is fixed at 109.47. To make the water molecule more rigid during the MD simulation, the SHAKE algorithm was used to constrain bond length and angle [28]. Long-range electrostatic interactions were computed using the Particle-Particle Particle-Mesh (PPPM) method [29]. The PPPM method is a distinct method for efficiently computing long-range electrostatic interactions in periodic systems [30]. It offers a complementary approach to the Ewald method, often demonstrating superior computational performance for large systems [31]. Non-bonding interactions, including both Van der Waals and electrostatic forces, were truncated at a cut-off distance of 12 Å. The interactions between Ti and O atoms were modeled using the 12-6 LJ potential, with parameters determined by the Lorentz-Berthelot mixing rule [32], as shown in Equ. 2.
@@
It should be noted that the interactions between Ti and H atoms are neglected in this study. This simplification based on the omission of LJ parameters for H atoms in the SPC/E water model is a well-established practice supported by extensive research and literature. The primary reasons for this simplification are computational efficiency and the focus to accurately capture the critical electrostatic interactions that determine the behavior of the water molecules [33, 34]. Additionally, because the bonds in the SPC/E model of water are rigid, the H atoms are restricted from freely rotating and forming independent interactions with Ti [26]. Therefore, the H atoms are primarily involved in H bonding with neighboring water molecules [33], further reducing their availability for direct interaction with Ti. It is worth mentioning that neglecting H-substrate interaction may underestimate the influence of wetting behavior on highly reactive (laser-treated) surfaces that can readily form H bonds with the H atoms of water. The LJ interaction potential parameters for Ti and water is found in Table 1.
@@
width = colspec = |Q[1,c,m]|Q[1,c,m]|Q[1,c,m]|Q[1,c,m]|Q[1,c,m]|, hlines, vlines, row1 = font=, rowodd = gray9, Surface (Surf) h (nm) w (nm) s (nm) p (nm)
@@
Figure 2: (a) A fence-like nanostructure on Ti surface with heights ($h$), widths ($w$), spacing ($s$), and the periodicity ($p$). The blue and red markers indicate non-irradiated and laser-irradiated regions, respectively, mimicking the rippled metasurfaces commonly produced by femtosecond laser processing. (b) Initial atomistic configuration of water nanodroplet on the flat Ti surface.
In addition, Ti-Ti interactions were calculated using the LJ potential. We initially tried LJ potentials from other sources [23, 24, 25, 36] including the Universal Force Field (UFF) [37]. The LJ potential parameters for Ti interactions used in this study are modified from the UFF in such a way that they can replicate the physical properties of Ti sufficient for Ti-water interactions that best aligns with existing calculations of water nanodroplets on materials [27]. The parameters have been validated against Density Functional Theory (DFT) calculations [38, 39] and experimental data [40, 41]. These parameters accurately reproduce the lattice parameters and independent elastic constants of Ti.
The nanodroplet corresponds to a sphere with a radius of 3.0 $nm$ and contains $3009$ atoms. Previous studies have consistently demonstrated that droplets comprising 2000 atoms (or more) consistently produce reliable results [42]. Temperature equilibration of the nanodroplet and the Ti substrate was performed separately. The nanodroplet was equilibrated using a time step of 1 $fs$ at 300 $K$ for approximately 1 $ns$ in the NVT ensemble. This duration was sufficient to achieve a stable nanodroplet configuration. To validate this configuration, the radial distribution functions (g(r)) obtained from our simulations were compared with the experimental data [35] and the SPC/E model [26], as shown in Fig. 1. The comparison indicates good agreement between our study, the experimental data, and the SPC/E model with slight deviations observed below the peaks of g(r). This agreement demonstrates that the SPC/E water model used in our MD simulations is appropriate and reliable.
Figure 3: Snapshots of the side and front views of the spreading process of a water nanodroplet on Ti surfaces with different roughness scales as a function of time. The roughness parameters used in the simulations are shown in Table 2: h (height), w (width), s (spacing) and period (p) vary across surfaces, with h ranging from 0.28 nm to 0.78 nm, w from 0.509 nm to 1.011 nm, s from 0.511 nm to 1.109 nm, and p from 0 to 2.02 nm respectively.
The flat Ti substrate corresponds to a block of dimension $14.5$$\times$ $19.5$$\times$ $1.4$ nm3 containing $22344$ atoms arranged in a Hexagonal Close-Packed (HCP) structure along [0-110] [10-10] [0001] directions. Subsequently, a nanometric pattern similar to laser-textured ones, having heights ($h$), widths ($w$), spacing ($s$), and a periodicity ($p$) as depicted in Fig. 2 (a), was introduced on Ti surfaces [43, 44, 45]. In our previous studies, it has been established that the structural parameters $h,w,s,$ and $p$ are adjustable based on the topography profile of the laser-patterned surfaces [46, 43]. The adjustable parameters used in the wetting studies are summarised in Table 2. For the sack of clarity and to facilitate the adjustment of these parameters, we use the roughness factor ($r$) and fractional area ($f$) to quantify them, as shown in Equ. 3 below.
@@
The bulk Ti substrate was equilibrated using the NVT ensemble at 300 $K$ for 1 $ns$ with a time step of 1 $fs$. In addition, the equilibrated spherical nanodroplet was positioned on the equilibrated Ti surface, as shown in Fig. 2 (b). In this study, the flat and rough Ti substrates were completely frozen during the MD simulation. The fixed substrate in MD simulations, particularly in wetting studies, is a well-documented practice aimed at a reduced computational cost [33], maintaining numerical stability [47] and ensuring an accurate representation of the physical system [48, 49]. Interestingly, simulations with a fixed surface compared to a flexible surface have shown a negligible effect on the CA [27].
In addition to the simulation set-up, the size of the simulation box was set to $14.5$$\times$ $19.5$$\times$ $15.4$ nm3. Periodic boundary conditions were applied in the x, y, and z directions during the simulation of the water nanodroplet on the Ti surfaces. The entire simulation was carried out for 1 ns in the NVT ensemble with a time step of 1 fs. Temperature control was achieved using the Nosé-Hoover thermostat to maintain a constant temperature of 300 K. The Velocity Verlet algorithm was used for numerical integration of the equations of motion. MD calculations were performed using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) software [50]. Visualization and analysis of the water nanodroplet on Ti surfaces were conducted using OVITO software [51].
@@ 3.1 Dynamic wetting behavior of Ti surfaces
To understand the fundamental process of nanodroplet spreading on the flat and nanometric Ti surfaces, the dynamics of the nanodroplet on the corresponding surfaces are examined. Thus, Fig. 3 shows the snapshots of the side and front views of the spreading process of a water nanodroplet on Ti surfaces with different roughness scales as a function of time. The droplet on the flat surface (Surf-1) is initially spherical at 0 ps, indicating minimal spreading. Between 30 and 50 ps, the droplet starts to spread gradually, with a noticeable decrease in height. From 200 to 1000 ps, further spreading occurs, and the droplet becomes more stable with a large base and minimal height.
@@
In the case of Surf-2, the droplet spreads and begins to interact with the rough surface by 30 ps. As time progresses, between 200 and 1000 ps, the droplet flattens out, forming a wide base that settles into the roughness grooves. This indicates a substantial degree of spreading. In Surf-3, the droplet spreading is restricted compared to Surf-2 and Surf-4. The narrow width and height of the grooves beneath the droplet in Surf-3 limit the extent to which the droplet can spread. The droplet conforms to these grooves, resulting in less lateral spreading compared to surfaces with larger grooves like Surf-2. This behavior is due to the greater groove density in Surf-3, as indicated by the equal distribution of w and s.
@@
Fig. 5 illustrates the dynamic CA of water nanodroplets as a function of time for Ti surfaces with varying roughness levels. The dynamic CA was determined by taking snapshots of the nanodroplets at regular intervals and measuring the advancing and receding CAs specifically in the Y-Z direction neglecting X-Z direction. The difference between the advancing and receding contact angles is represented by the error bars. Initially, the droplets exhibit minimal spreading on all surfaces, with the CA at 30 ps ranging from approximately $115^{\circ}$ for the flat surface (Surf-1) to around $133^{\circ}$ for Surf-4. The CA decreases over time, signifying the spreading of the droplets, with the most rapid decrease occurring within the initial 50 ps. This decrease is due to the dynamic rearrangement of water molecules at the liquid-Ti interface, driven by intermolecular forces and the electrostatic interaction between the oxygen atom of the water molecule and the Ti atoms. This molecular rearrangement lowers the interfacial energy, resulting in a decrease in the CA, which leads to a more stable and spread-out droplet configuration. Notably, Surf-5, with the highest roughness, consistently exhibits a higher CA compared to the other surfaces, indicating that increased roughness hinders wetting and promotes hydrophobicity due to the trapping of air pockets beneath the droplet.
@@
Thus, in general, the CA increases as roughness rises shifting the wettability of Ti surfaces from hydrophilic (Surf-1) to hydrophobic state for other surfaces. These observations for Surf-2, 3, and 4 align with the main assumption of the Wenzel model. The Cassie-Baxter model somewhat better describes Surf-5, where air pockets exist between the droplet and the rough Ti surface, generally leading to higher CAs and less wetting. Interestingly, despite small sizes, the initial rapid decrease in CA and an increase in spreading radius is similar to the dynamic wetting theories and previous simulation studies, where the droplet adjusts its shape quickly upon initial contact to minimize the surface energy. Hence, the simulation model replicates key features of nanotextured surfaces created by femtosecond laser processing, using simplified nanostructures for better control and precision [4, 52, 53]. While experimentally obtained laser-induced textures have some level of complexity and irregularity, the geometric parameters (e.g., height, width, spacing, and period) were based on experimental observations to ensure physical relevance . This approach allows us to focus on essential nanodroplet-surface interactions while maintaining computational efficiency and providing insights applicable to real-world conditions [43].
It is important to mention that the nanodroplet undergoes significant vibration on all the surfaces from 50 ps to 1000 ps, as seen in Fig. 3, 4 and 5. The damping of droplet vibrations on rough surfaces is driven by several key mechanisms. As the surface roughness increases, the contact area between the droplet and surface also increases, leading to higher energy dissipation at the liquid-solid interface [54, 55, 56]. Surfaces with deeper grooves (as observed in Surf-5) tend to trap air pockets beneath the droplet, which creates additional resistance against droplet movement. This increased resistance helps to dampen the vibrations more effectively compared to smoother surfaces. Conversely, surfaces like Surf-2 with moderate roughness facilitate higher spreading but lower vibration damping, as the droplet conforms more closely to the surface grooves without substantial energy dissipation from trapped air pockets.
The vibrations are also influenced by the droplet’s interaction with nanoscale surface features, where molecular-level forces, including van der Waals forces and electrostatic interactions, dominate [42]. The enhanced interaction between water molecules and the Ti surface atoms at nanometer roughness increases energy transfer from the droplet to the surface, contributing to the observed vibration damping [57]. The findings have significant implications for practical applications, particularly in the biomedical field where Ti surfaces are frequently used for implants [11]. The vibration damping effect observed with nanostructured surfaces improves the implant’s performance by ensuring stable liquid interaction, thus improving biofluid adhesion and minimizing the risk of bacterial attachment, which is important for preventing infections.